Solve algebraically, using one variable. The sum of two positive integers is 25. The larger integer is five less than four times the smaller one. What are the two integers?
Answer by nerdybill(7384)

Solve algebraically, using one variable. The sum of two positive integers is 25. The larger integer is five less than four times the smaller one. What are the two integers?
.
Let x = smaller of the two numbers
then because "The sum of two positive integers is 25"
25-x = larger of two numbers
.
25-x = 4x - 5
25 = 5x - 5
30 = 5x
6 = x (smaller number)
.
Larger number:
25-x = 25-6 = 19 (larger number)
.
The two integers are 6 and 19.